# Multi-class Temporal Logic Neural Networks

### v1

@@ Abstract
Time-series data can represent the behaviors of autonomous systems, such as drones and self-driving cars. The problem of binary and multi-class classification has received a lot of attention in this field. Neural networks represent a popular approach to classifying data; However, they lack interpretability, which poses a significant challenge in extracting meaningful information from them. Signal Temporal Logic (STL) is a formalism to describe the properties of timed behaviors. We propose a method that combines all of the above: neural networks that represent STL specifications for multi-class classification of time-series data. We offer two key contributions: 1) We introduce a notion of margin for multi-class classification, and 2) we introduce the use of STL-based attributes for enhancing the interpretability of the results. We evaluate our method on two datasets and compare with state-of-the-art baselines.
@@ I INTRODUCTION
Multi-class classification is an essential problem in Machine Learning (ML). In this paper, we are interested in classifying behaviors from time-series data: given a set of signals sampled over a period of time, extract features that allow the signal to be assigned to one of different known categories of interest. Many ML algorithms have been proposed to solve this type of multi-class classification problem; however, most of these produce models that are hard to interpret by a human, and it is, in general, difficult to identify the attributes (features) of a signal that distinguish one category from the others.
@@
Recently, temporal logic tools have attracted lots of attention due to the rich expressiveness and natural interpretability of temporal logic formulae. Signal Temporal Logic (STL) is an expressive formal language to describe the properties of time series. The quantitative satisfaction of STL formulae provides a formal way to assess how close a signal is to satisfy a given specification; the specification itself can also be expressed in a language that can be easily understood by humans. There have been encouraging results that return STL formulae for classification. However, most of the existing work only considers binary classification tasks; this is probably due to the fact that the semantics of STL are tightly related to boolean logic, hence offering a natural way to distinguish only between two (positive versus negative) classes. Few studies have been done to introduce STL technique to a multi-class classifier.
@@
Attribute-based classification is a classification method intuited by how humans distinguish objects[1]. Humans detect different objects from a high-level description comprising arbitrary semantic attributes. The authors in [1] propose a system that can detect a list of attributes that can transcend between the classes. The attribute-based classification is statistically efficient since the knowledge about attributes can be obtained from different classes. New classes that are not observed during training can be detected based on their attribute representation without the necessity for an additional training phase, which is also known as zero-shot learning. However, most attribute-based classification methods focus on classifying images instead of time-series data[2, 3, 4].
@@
Various techniques have been proposed to solve multi-class classification problems[5]. One of the traditional approaches is to convert the multi-class problems to multiple binary classification problems, which can be solved using existing binary classification techniques, such as Support Vector Machine (SVM) [6]. Others solve the problem by using algorithms that can directly handle multiple classes. For instance, one of the simplest methods is $k$-Nearest Neighbors, which separates the data into $k$ groups based on the distance to $k$ class of training examples [7]. A more advanced technique that produces more compact models is Decision Trees (DT), which splits the data using sequential decisions across nodes arranged in a tree, and where the leafs of the tree represent class decisions [8]. Neural networks designate each neuron in the output layer as the identification for one class[9]. However, these techniques produce models that are either hard to interpret or too complex to combine with the idea of attributes.
Efforts have been made to incorporate temporal logic tools as a way to improve the interpretability of the classification models. Existing work includes combining the neural network with weighted signal temporal logic[10, 11, 12, 13], extending SVMs to Support-Vector Machine-Signal Temporal Logic (SVM-STL)[14], building a decision tree that can be represented as signal temporal logic formulae[15]. These methods, however, focus only on binary classification based on the signs of the quantitative satisfaction of the signal temporal logic formulae.
@@
We are aware of two previous works for multi-class classification with DTs STL specifications [16, 14]. Both methods learn the parameters of Parametric Signal Temporal Logic (PSTL). Nevertheless, the DT method is not able to employ attribute learning since a signal ends up at only a single node of DT but may satisfy multiple attributes simultaneously. The efficiency of the decision tree method is restricted by its depth and the sample size of the dataset. Moreover, they only consider the signs (satisfaction or violation) of the STL formulae and discard the quantitative semantics.
@@
Contributions: First, we propose a neural network that can be translated to STL formulae for specifying the attributes of signals. Second, we propose a novel margin-based loss function that leverages the quantitative semantics of the STL formulae and that explicitly improves the separation between the data. Third, we evaluate our approach on two datasets and highlight the performance of our approach compared to existing methods.
@@ II-A Error-Correcting Output Codes
The ECOC approach considers the label of each class as a unique codeword, which is an arbitrary binary string. The codewords are mapped to a binary coding matrix $E$ with $c$ rows and $n$ columns, where $c$ is the number of classes and $n$ is the number of codeword bits. The length of the codeword, i.e., the number of columns of the coding matrix, needs to be sufficient to represent c different labels. The codeword for each class is decided before the training; in general, it is selected to maximize its separation distance from the other codewords. The classifier $f$ is defined by aggregating the output of $n$ binary classifiers $f_{1},f_{2},\ldots,f_{n}$, each corresponding to one column of the ECOC matrix.
The coding matrix $E$ takes the values $\{1,-1\}$ to integrate with the STL inference method. Table I shows an example of a coding matrix, where $+$ represents $+1$ and -- represents $-1$. During the training process, the grounded-truth output of the $i$-th binary classifier $f_{i}$ is specified by the $i$-th bit of these codewords. For example, $f_{2}(x)=1$ if $x$ belongs to class $1$ or class $4$, $f_{2}(x)=-1$ if $x$ belongs to class $2$ or class $3$. The prediction of the classifier $f$ is based on the distance between the codeword obtained by $f_{1},\ldots,f_{n}$ and the codewords in the matrix. In our context, each classifier is correlated to an attribute.
$f_{1}$
$f_{2}$
$f_{3}$
Class 1
Class 2
Class 3
Class 4
@@ II-B Signal Temporal Logic
Signal Temporal Logic (STL) is a tool for specifying the temporal properties of time-series data. We denote a discrete-time, $d$-dimensional, time-series signal $s$ as a finite sequence $s(1),s(2),\ldots$, where $s(\tau)\in d$ is the value of signal $s$ at time $\tau$. The syntax of STL formulae is defined recursively as [18]:
@@ II-C NN-TLI
The Neural Network (NN) architecture for Temporal Logic Inference (TLI), called NN-TLI, is a method for binary classification[10]. The neural network consists of three layers: a predicate layer, a temporal operator layer and a Boolean operator layer. The network can be translated to a level-1 STL formula $\phi$ in the disjunctive normal form (DNF):
$\phi=(\varphi_{11}\land\varphi_{12}\land\ldots\land\varphi_{1q_{1}})\lor\ldots\lor(\varphi_{p1}\land\varphi_{p2}\land\ldots\land\varphi_{pq_{n}}),$ (3)
such that for a signal $s$, if $s\models\phi$, it is predicted as positive label; if $s\not\models\phi$, it is predicted as negative label.
To construct the STL formula $\phi$ in the form of (3), the network first learns $m$ candidate sub-formulae $\varphi$, each corresponding to a neuron of the temporal operator layer. In the Boolean operator layer, the network formulates the final STL formula using a conjunction-disjunction matrix $M$. The conjunction-disjunction matrix $M$ is a binary matrix defined as:
$M=\begin{bmatrix}c_{11}&c_{12}&\ldots&c_{1q_{m}}\\ c_{21}&c_{22}&\ldots&c_{2q_{m}}\\ \vdots&\vdots&\ddots&\vdots\\ c_{p_{m}1}&c_{p_{m}2}&\ldots&c_{p_{m}q_{m}}\\ \end{bmatrix}\in\{0,1\}^{p_{m}\times q_{m}}$ (4)
where $p_{m}>p$, $q_{m}>\max\{q_{1},\ldots,q_{n}\}$. The network takes conjunction inside each row followed by disjunction between the rows to construct the STL formula in the form of (3). For example, for $3$ candidate subformulae $\varphi_{1}$, $\varphi_{2}$, $\varphi_{3}$ and $M=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&1&1\end{bmatrix}$, the learned STL formula is $\phi=\varphi_{1}\lor(\varphi_{2}\land\varphi_{3})$.
@@ III Problem Formulation
We aim to capture the temporal attributes of signals, learning by STL formulae, and use these attributes for multi-class classification.
In attribute-based classification [1], the attribute representation $\mathbf{a}^{y}=\begin{bmatrix}a_{1}^{y},\ldots,a_{n}^{y}\end{bmatrix}$ for class $y$ is a fixed-length binary vector. We map the binary attribute representation to an ECOC table and consider the classifier for each attribute as an STL formula. The output of an STL formula is either a positive or non-positive value that is mapped to an attribute label. We can then use the prediction method of ECOC approach to predict the class of a time-series signal.
We define a multi-class classification task with mutually exclusive classes, i.e., where each signal belongs to one and only one class. Consider a dataset with $N$ discrete-time signals $\mathcal{S}=\{(s^{i},\mathbf{a}^{i},y^{i})\}_{i=1}^{N}$, where $s^{i}$ is the $i$-th signal, $\mathbf{a}^{i}\in\{-1,+1\}^{n}$ is the attribute label of the signal, $y^{i}\in\mathcal{C}$ is the class of the signal, $\mathcal{C}=\{1,\ldots,c\}$ is the set of classes, $\lvert\mathcal{C}\rvert=c$. Our goal is to learn $n$ STL formulae $\phi_{1},\phi_{2},\ldots,\phi_{n}$ in the form of (3) such that these $n$ STL formulae can be used as the binary classifiers $f^{\prime}s$ explained in Section II-A and decide the class it belongs to. Based on the relation between attributes and classes, we introduce two forms of the problem:
@@
$\forall(s,\mathbf{a},y)\in\mathcal{S},(s\models\phi_{i},\forall a_{i}=1)\land(s\not\models\phi_{j},\forall a_{j}=-1)$ (5)
@@
$\forall(s,\mathbf{a},y)\in\mathcal{S},(s\models\psi_{y},y\in\mathcal{C})\land(s\not\models\psi_{y^{\prime}},\forall y^{\prime}\in\mathcal{C}\backslash y)$ (6)
@@
Note that if a signal satisfies the class decoder formula of class $y$, it automatically satisfies the attributes specified by the attribute label $\mathbf{a}^{y}$, and vice versa; hence, these two problem formulations are semantically the same. However, they lead to different forms for the loss function design.
@@ IV Method
We propose a method for multi-class classification of time-series signals. Our approach is based on three components. First, we modify the NN-TLI binary classification architecture to a multiclass setting. Second, we introduce a novel notion of margin based on the multi-dimensional embedding produced by the formulae and define a loss function that uses this notion. Finally, we apply loss-based decoding based on the error-correcting output code (ECOC) for the final prediction.
@@ IV-A Multi-class NN-TLI
Our proposed multi-class NN-TLI architecture is a generalization of the binary NN-TLI architecture and uses a new conjunction-disjunction matrix $M_{m}$ for multi-class problems, which we call the multi-conjunction-disjunction matrix. For a $c$-class classification problem, we define $M_{m}$ as:
$M_{m}=\begin{bmatrix}M_{1}^{T},M_{2}^{T},\ldots,M_{n}^{T}\end{bmatrix}^{T}\in\{0,1\}^{n\times p_{m}\times q_{m}}$ (8)
where $M_{i}$ is a conjunction-disjunction matrix defined in (4). Each matrix $M_{i}$ defines an attribute formula $\phi_{i}$, which is a Boolean combination of a subset of the sub-formulae pool. These attributes formulae are learned simultaneously thus can increase training efficiency and accuracy compared to learning $n$ independent neural networks.
@@ IV-B Multi-class margin and margin-based Loss
In the rest of this paper, we denote $r_{ak}$ as the quantitative robustness of the attribute formula $\phi_{k}$ given a sample; $E$ as the ECOC coding matrix; $E(j,\cdot)$ as the $j$-th row of the matrix $E$; $E(\cdot,k)$ as the $k$-th column of the matrix $E$; $E(j,k)$ as the $k$-th bit of the $j$-th row of the matrix $E$. For the class-based classification (Problem 2), we define a class vector $\mathbf{r}_{y}(s)$:
@@
$r(s,\psi_{j})=\min_{k=1,\ldots,n}E(j,k)r_{ak},\forall j\in\mathcal{C}$ (11)
$r_{yj}$ is the quantitative robustness of the class formula $\psi_{j}$ given a sample. The output vector $\mathbf{r}(s)$ of the multi-class NN-TLI is either an attribute vector or a class vector.
Inspired by the concept of margin in SVM [19], we introduce a novel notion of margin for multi-class STL inference. In traditional binary SVM classifiers, the margin is defined by the minimum distance between a training sample of either class and the separating hyperplane. The training algorithm then aims to find a hyperplane that maximizes this margin. In our multi-class NN-TLI, instead, we interpret each formula as a hyperplane that is aligned with the axis. Differently from SVM, the training algorithm does not change the position of this separating hyperplane but changes the position of the output vector $\mathbf{r}$ for each sample. We therefore propose a new notion of margin for the multi-class STL classification.
@@
we then introduce the margin for multi-class STL as follows:
@@
[…]
@@
[…]
(a) 3D output space
(b) Project 1D
Figure 1: Example of a 3D output space.
@@
Consider a three-class classification task with a class set $\{1,2,3\}$ and a class-based classification. The output space created by multi-class NN-TLI is $3$. The network maps the training samples to three-dimensional points, as shown in Figure 1a. We can project the points to one of the axes, e.g., the $x$-axis subspace shown in Figure 1b. From the problem statement, the points with positive $x$-coordinate are considered to satisfy $\psi_{1}$, and those with negative $x$-coordinate violate $\psi_{1}$. The margin of $\psi_{1}$ is $m_{y1}=\min\{m_{p},m_{n}\}$.
@@
$L_{a}(\mathbf{r}_{a},E)$ $=\sum_{k=1}^{n}\sum_{i=1}^{N}\Big{(}\text{ReLU}\big{(}m_{ak}-E(y^{i},k)r_{ak}(s^{i})\big{)}-\delta m_{ak}\Big{)},$ (15)
while for the class-based classification problem is
$L_{y}(\mathbf{r}_{y},C)$ $=\sum_{j=1}^{c}\sum_{i=1}^{N}\Big{(}\text{ReLU}\big{(}m_{yj}-C(y^{i},j)r_{yj}(s^{i})\big{)}-\delta m_{yj}\Big{)},$ (16)

### v2

@@ Abstract
Time-series data can represent the behaviors of autonomous systems, such as drones and self-driving cars. The task of binary and multi-class classification for time-series data has become a prominent area of research. Neural networks represent a popular approach to classifying data; However, they lack interpretability, which poses a significant challenge in extracting meaningful information from them. Signal Temporal Logic (STL) is a formalism that describes the properties of timed behaviors. We propose a method that combines all of the above: neural networks that represent STL specifications for multi-class classification of time-series data. We offer two key contributions: 1) We introduce a notion of margin for multi-class classification, and 2) we introduce STL-based attributes for enhancing the interpretability of the results. We evaluate our method on two datasets and compare it with state-of-the-art baselines.
@@ I INTRODUCTION
Multi-class classification is an essential problem in Machine Learning (ML). In this paper, we are interested in classifying behaviors from time-series data: given a set of signals sampled over a period of time, we extract features that allow the signal to be assigned to one of the different known categories of interest. Many ML algorithms have been proposed to solve this type of multi-class classification problem; however, most of these algorithms produce models that are hard to interpret by a human, and it is, in general, difficult to identify the attributes (features) of a signal that distinguish one category from the others. Recently, temporal logic tools have attracted lots of attention due to the rich expressiveness and natural interpretability of temporal logic formulae. Signal Temporal Logic (STL) is an expressive formal language to describe the properties of time-series data. The quantitative satisfaction of STL formulae provides a formal way to assess how close a signal is to satisfy a given specification; the specification itself can also be expressed in a language understandable to humans. There have been encouraging results that return STL formulae for classification. However, most of the existing work only considers binary classification tasks [1, 2, 3]; this is probably due to the fact that the semantics of STL are tightly related to boolean logic, hence offering a natural way to distinguish only between two (positive versus negative) classes. Few studies have been done to introduce the STL technique to a multi-class classifier.
@@
Attribute-based classification is a method intuited by how humans distinguish objects[4]. Humans detect different objects from a high-level description comprising arbitrary semantic attributes. The authors in [4] propose a system that can detect a list of attributes transcending between the classes. The attribute-based classification is statistically efficient since the knowledge about attributes is obtained from different classes. New classes not observed during training can be detected based on their attribute representation without the necessity for an additional training phase, which is also known as zero-shot learning. In the domain of autonomous systems, such as robots or drones, their trajectories typically involve the execution of a series of tasks. Tasks often overlap or intersect within a group of autonomous entities to facilitate efficient coordination. The attribute-based classification method categorizes autonomous system trajectories by discerning and learning individual task specifications. However, most attribute-based classification methods focus on classifying images instead of time-series data[5, 6, 7].
@@
Various techniques have been proposed to solve multi-class classification problems[8]. One of the traditional approaches is to convert the multi-class task to multiple binary classification tasks, which can be solved using existing binary classification techniques, such as Support Vector Machine (SVM) [9]. Others solve the problem by using algorithms that can directly handle multiple classes. For instance, one of the simplest methods is $k$-Nearest Neighbors, which separates the data into $k$ groups based on the distance to $k$ class of training examples [10]. A more advanced technique that produces more compact models is Decision Trees (DT), which splits the data using sequential decisions across nodes arranged in a tree, and where the leaves of the tree represent class decisions [11]. Neural networks designate each neuron in the output layer as the identification for one class[12]. However, these techniques produce models that are either hard to interpret or too complex to combine with the idea of attributes.
@@
Efforts have been made to incorporate temporal logic tools to improve the interpretability of the classification models. Existing work includes combining the neural network with weighted signal temporal logic[13, 14, 15, 16], extending SVMs to Support Vector Machine-Signal Temporal Logic (SVM-STL)[17], building a decision tree that can be represented as signal temporal logic formulae[3]. These methods, however, focus only on binary classification based on the signs of the quantitative satisfaction of the signal temporal logic formulae. We are aware of two previous works for multi-class classification with DTs STL specifications [18, 17]. Both methods learn the parameters of Parametric Signal Temporal Logic (PSTL). Nevertheless, the DT method is not able to employ attribute learning since a signal ends up at only a single node of DT but may satisfy multiple attributes simultaneously. The efficiency of the decision tree method is restricted by its depth and the sample size of the dataset. Moreover, they only consider the signs (satisfaction or violation) of the STL formulae and discard the quantitative semantics.
@@
Contributions: First, we propose a neural network that can be translated to STL formulae for specifying the attributes of signals. Second, we propose a novel margin-based loss function that leverages the quantitative semantics of the STL formulae and explicitly improves the separation between the data. Third, we evaluate our approach on two datasets and highlight the performance of our approach compared to existing methods.
@@ II-A Error-Correcting Output Codes
The ECOC approach considers the label of each class as a unique codeword, i.e., an arbitrary binary string. The codewords are mapped to a binary coding matrix $E$ with $c$ rows and $n$ columns, where $c$ is the number of classes and $n$ is the number of codeword bits. The length of the codeword, i.e., the number of columns of the coding matrix, must be sufficient to represent $c$ unique labels. The codeword for each class is decided before the training; in general, it is selected to maximize its separation distance from the other codewords. The classifier $f$ is defined by aggregating the output of $n$ binary classifiers $f_{1},f_{2},\ldots,f_{n}$, each corresponding to one column of the coding matrix.
The coding matrix $E$ takes the values $\{1,-1\}$ to integrate with the STL inference method. Table I shows an example of a coding matrix, where $+$ represents $+1$ and -- represents $-1$. During the training process, the grounded-truth output of the $i$-th binary classifier $f_{i}$ is specified by the $i$-th bit of these codewords. For example, $f_{2}(x)=1$ if $x$ belongs to class $1$ or class $4$, $f_{2}(x)=-1$ if $x$ belongs to class $2$ or class $3$. The prediction of the classifier $f$ is based on the distance between the codeword obtained by $f_{1},\ldots,f_{n}$ and the codewords in the matrix. In our context, each classifier is correlated to an attribute specified by an STL formula.
$f_{1}$ $f_{2}$ $f_{3}$
Class 1 - + -
Class 2 - - +
Class 3 + - -
Class 4 - + +
@@ II-B Signal Temporal Logic
Signal Temporal Logic (STL) is a tool for specifying the spatial-temporal properties of time-series data. We denote a discrete-time, $d$-dimensional, time-series signal $s$ as a finite sequence $s(1),s(2),\ldots,s(l)$, where $s(\tau)\in d$ is the value of signal $s$ at time $\tau$, $l$ is the length of signal. The syntax of STL formulae is defined recursively as [20]:
@@ II-C NN-TLI
The Neural Network (NN) architecture for Temporal Logic Inference (TLI), called NN-TLI [13], is a method for binary classification. The NN-TLI learns an STL formula to classify desired and undesired behaviors as satisfying and violating the formula, respectively. The neural network consists of three layers: a predicate layer, a temporal layer, and a Boolean layer. The predicate layer learns $q_{m}$ predicates, where $q_{m}$ is the number of neurons in this layer. The temporal layer learns the time interval corresponding to these $q_{m}$ predicates. In the Boolean layer, the network formulates the final STL formula using a conjunction-disjunction matrix $M$. The conjunction-disjunction matrix $M$ is a binary matrix defined as:
$M=\begin{bmatrix}c_{11}&c_{12}&\ldots&c_{1q_{m}}\\ c_{21}&c_{22}&\ldots&c_{2q_{m}}\\ \vdots&\vdots&\ddots&\vdots\\ c_{p_{m}1}&c_{p_{m}2}&\ldots&c_{p_{m}q_{m}}\\ \end{bmatrix}\in\{0,1\}^{p_{m}\times q_{m}},$ (3)
where $q_{m}$ and $p_{m}$ are hyperparameters of NN-TLI. The network learns the parameters of matrix $M$; it takes conjunction inside each row, followed by a disjunction between the rows based on the elements of the matrix. The learned STL formula $\phi$ is a level-1 STL formula in the disjunctive normal form (DNF):
$\phi=(\varphi_{11}\land\varphi_{12}\land\ldots\land\varphi_{1q_{1}})\lor(\varphi_{21}\land\varphi_{22}\land\ldots\land\varphi_{2q_{2}})\lor\\ \ldots\lor(\varphi_{p1}\land\varphi_{p2}\land\ldots\land\varphi_{pq_{n}}),$ (4)
note that $p_{m}>p$, $q_{m}\geq\max\{q_{1},\ldots,q_{n}\}$. The conjunction-disjunction matrix only specifies the maximum complexity of the formula, which provides the flexibility to acquire a concise STL formula by learning elements $c_{ij}$’s of matrix $M$. For example, if the predicate layer and temporal layer learn $3$ subformulae $\varphi_{1}$, $\varphi_{2}$, $\varphi_{3}$, and the learned conjunction-disjunction matrix is a $3\times 3$ matrix $M=\begin{bmatrix}1&0&0\\ 0&0&0\\ 0&1&1\end{bmatrix}$, the final STL formula is $\phi=\varphi_{1}\lor(\varphi_{2}\land\varphi_{3})$. For a signal $s$, if $s\models\phi$, it is predicted as the positive label; if $s\not\models\phi$, it is predicted as the negative label. Details of the learning process can be found in [13].
@@ III Problem Formulation
We aim to capture the spatial-temporal attributes of signals, learning by STL formulae, and use these attributes for multi-class classification.
In attribute-based classification [4], the attribute representation $\mathbf{a}^{y}=\begin{bmatrix}a_{1}^{y},\ldots,a_{n}^{y}\end{bmatrix}$ for class $y$ is a fixed-length binary vector. We map the binary attribute representation to an ECOC table and consider the classifier for each attribute as an STL formula. The output of an STL formula is either a positive or non-positive value mapped to an element of $\mathbf{a}^{y}$. We can then use the prediction method of the ECOC approach to predict the class of a time-series signal.
We define a multi-class classification task with mutually exclusive classes, i.e., where each signal belongs to one and only one class. Consider a dataset with $N$ discrete-time signals $\mathcal{S}=\{(s^{i},\mathbf{a}^{i},y^{i})\}_{i=1}^{N}$, where $s^{i}\in d\times l$ is the $i$-th signal, $l$ is the length of the signal; $\mathbf{a}^{i}\in\{-1,+1\}^{n}$ is the attribute label of the signal; $y^{i}\in\mathcal{C}$ is the class of the signal; $\mathcal{C}=\{1,\ldots,c\}$ is the set of classes, $\lvert\mathcal{C}\rvert=c$. Our goal is to learn $n$ STL formulae $\phi_{1},\phi_{2},\ldots,\phi_{n}$ in the form of (4) such that these $n$ STL formulae can be used as the binary classifiers $f^{\prime}s$ explained in Section II-A and decide the class it belongs to. Based on the relation between attributes and classes, we introduce two forms of the problem:
@@
$\forall(s,\mathbf{a},y)\in\mathcal{S},(s\models\phi_{i},\forall a_{i}=1)\land(s\not\models\phi_{j},\forall a_{j}=-1).$ (5)
@@
$\forall(s,\mathbf{a},y)\in\mathcal{S},(s\models\psi_{y},y\in\mathcal{C})\land(s\not\models\psi_{y^{\prime}},\forall y^{\prime}\in\mathcal{C}\backslash y),$ (6)
@@
Note that if a signal satisfies $\psi_{y}$, it automatically satisfies the attributes specified by the attribute label $\mathbf{a}^{y}$, and vice versa; hence, these two problem formulations are semantically the same. However, they lead to different forms for the loss function design.
@@ IV Method
We propose a method for multi-class classification of time-series signals. Our approach is based on three components. First, we modify the NN-TLI binary classification architecture to a multiclass setting. Second, we introduce a novel notion of margin based on the multi-dimensional embedding produced by the formulae and define a loss function that uses this notion. Finally, we apply loss-based decoding based on the error-correcting output codes (ECOC) for the prediction.
@@ IV-A Multi-class NN-TLI
Our proposed multi-class NN-TLI architecture is a generalization of the binary NN-TLI architecture. It uses a new conjunction-disjunction matrix $M_{m}$ for multi-class problems, which we call the multi-conjunction-disjunction matrix. For a $c$-class classification problem, we define $M_{m}$ as:
$M_{m}=\begin{bmatrix}M_{1}\\ \vdots\\ M_{n}\end{bmatrix}\in\{0,1\}^{n\times p_{m}\times q_{m}},$ (8)
where $M_{i}$ is a conjunction-disjunction matrix defined in (3). Each matrix $M_{i}$ defines an STL formula $\phi_{i}$, a Boolean combination of a subset of the subformulae pool. These attributes formulae are learned simultaneously thus can increase training efficiency and accuracy compared to learning $n$ independent neural networks.
@@ IV-B Multi-class Margin and Margin-based Loss
In the rest of this paper, we denote $r_{ak}$ as the robustness of $\phi_{k}$ given a sample; $E$ as the ECOC coding matrix; $E(j,\cdot)$ as the $j$-th row of the matrix $E$; $E(\cdot,k)$ as the $k$-th column of the matrix $E$; $E(j,k)$ as the $k$-th bit of the $j$-th row of the matrix $E$. For the class-based classification (Problem 2), we define a class vector $\mathbf{r}_{y}(s)$:
@@
$r(s,\psi_{j})=\min_{k=1,\ldots,n}E(j,k)r_{ak},\forall j\in\mathcal{C},$ (11)
$r_{yj}$ is the robustness of $\psi_{j}$ given a sample. The output vector $\mathbf{r}(s)$ of the multi-class NN-TLI is either an attribute vector or a class vector.
Inspired by the concept of margin in SVM [21], we introduce a novel notion of margin for multi-class STL inference. In traditional binary SVM classifiers, the margin is defined by the minimum distance between a training sample of either class and the separating hyperplane. The training algorithm then aims to find a hyperplane that maximizes this margin. In our multi-class NN-TLI, instead, we interpret each formula as a hyperplane aligned with the axis. Unlike SVM, the training algorithm does not change the position of this separating hyperplane but changes the position of the output vector $\mathbf{r}$ for each sample. Therefore, we propose a new notion of margin for the multi-class STL classification.
@@
we then introduce the multi-class STL margin as follows:
@@
$\begin{gathered}m_{ak}=\min_{i}\{\text{ReLU}(E(y^{i},k)r_{ak}(s^{i}))\},\forall k\in\{1,\ldots,n\}\\ \text{ReLU}(x)=\begin{cases}x,&\textrm{if }x>0,\\ 0,&\textrm{otherwise}.\end{cases}\end{gathered}$ (13)
@@
$m_{yj}=\min_{i}\{\text{ReLU}(C(y^{i},j)r_{yj}(s^{i}))\},\forall j\in\{1,\ldots,c\}.$ (14)
(a) The three-dimensional output space
(b) The projection onto one-dimensional subspace
Figure 1: Example of a three-dimensional output space.
@@
Consider a three-class classification task with a class set $\{1,2,3\}$ and a class-based classification. The output space created by multi-class NN-TLI is $3$. The network maps the training samples to three-dimensional points, as shown in Figure 1a. We can project the points to one of the axes, e.g., the $x$-axis subspace shown in Figure 1b. From the problem statement, the points with positive $x$-coordinate satisfy $\psi_{1}$, and those with negative $x$-coordinate violate $\psi_{1}$. The margin of $\psi_{1}$ is $m_{y1}=\min\{m_{p},m_{n}\}$.
@@
$L_{a}(\mathbf{r}_{a},E)=\sum_{k=1}^{n}\sum_{i=1}^{N}\Big{(}\text{ReLU}\big{(}m_{ak}-E(y^{i},k)r_{ak}(s^{i})\big{)}\\ -\delta m_{ak}\Big{)},$ (15)
while the loss function for the class-based classification problem is
$L_{y}(\mathbf{r}_{y},C)=\sum_{j=1}^{c}\sum_{i=1}^{N}\Big{(}\text{ReLU}\big{(}m_{yj}-C(y^{i},j)r_{yj}(s^{i})\big{)}\\ -\delta m_{yj}\Big{)},$ (16)